OpenStudy (anonymous):

In a right isosceles triangle, the measure of line AB is 4. What is the measure of line AC? 2√(2) 4 16 4√(2) 32

OpenStudy (anonymous):

Is m(<A) = 90

OpenStudy (anonymous):

depends on order of A, B and C. if AB is leg, then AC = 4sqrt(2) if AB is hypotenuse, then AC = 2sqrt(2)

OpenStudy (anonymous):

Question must be more specific..

OpenStudy (anonymous):

AB can't be hypotenuse..

OpenStudy (anonymous):

why

OpenStudy (anonymous):

If AB is hypotenuse then BC will also 4 because it is isosceles.. And as this is right triangle hypotenuse must be greater of the other two..

OpenStudy (anonymous):

i attached a diagram

OpenStudy (anonymous):

Yep angle B is right angle and AC is hypotenuse..

OpenStudy (anonymous):

so what is is the measure of the hypotenuse?

OpenStudy (anonymous):

Use pythagorean theorem..

OpenStudy (anonymous):

\[AC^2 = AB^2 + BC^2\]

OpenStudy (anonymous):

Find AC from here... AB = BC = 4 here..

OpenStudy (anonymous):

would it be 16?

OpenStudy (anonymous):

Don't guess.. Use the above formula to find AC..

OpenStudy (anonymous):

\[AC^2 = (4)^2 + (4)^2 = ??\]

OpenStudy (anonymous):

32

OpenStudy (anonymous):

So: \[AC^2 = 32\] Take the Square root both the sides to find AC now..

OpenStudy (anonymous):

4√(2)

OpenStudy (anonymous):

Yep...